### v1

# The ELM Neuron: an Efficient and Expressive

Abstract
Traditional large-scale neuroscience models and machine learning utilize simplified models of individual neurons, relying on collective activity and properly adjusted connections to perform complex computations. However, each biological cortical neuron is inherently a sophisticated computational device, as corroborated in a recent study where it took a deep artificial neural network with millions of parameters to replicate the input-output relationship of a detailed biophysical model of a cortical pyramidal neuron. We question the necessity for these many parameters and introduce the Expressive Leaky Memory (ELM) neuron, a biologically inspired, computationally expressive, yet efficient model of a cortical neuron. Remarkably, our ELM neuron requires only 8K trainable parameters to match the aforementioned input-output relationship accurately. We find that an accurate model necessitates multiple memory-like hidden states and intricate nonlinear synaptic integration. To assess the computational ramifications of this design, we evaluate the ELM neuron on various tasks with demanding temporal structures, including a sequential version of the CIFAR-10 classification task, the challenging Pathfinder-X task, and a new dataset based on the Spiking Heidelberg Digits dataset. Our ELM neuron outperforms most transformer-based models on the Pathfinder-X task with $77\%$ accuracy, demonstrates competitive performance on Sequential CIFAR-10, and superior performance compared to classic LSTM models on the variant of the Spiking Heidelberg Digits dataset. These findings indicate a potential for biologically motivated, computationally efficient neuronal models to enhance performance in challenging machine learning tasks.
1 Introduction
The human brain has impressive computational capabilities, yet the precise mechanisms underpinning them remain largely undetermined. Two complementary directions are pursued in search of mechanisms for brain computations. On the one hand, many researchers investigate how these capabilities could arise from the collective activity of neurons connected into a complex network structure Maass (1997); Gerstner and Kistler (2002); Grüning and Bohte (2014), where individual neurons might be as basic as leaky integrators or ReLU neurons. On the other hand, it has been proposed that the intrinsic computational power possessed by individual neurons Koch (1997); Koch and Segev (2000); Silver (2010) contributes a significant part to the computations.
Even though most work focuses on the former hypothesis, an increasing amount of evidence indicates that cortical neurons are remarkably sophisticated Silver (2010); Almog and Korngreen (2016); Gidon et al. (2020); Larkum (2022), even comparable to expressive multilayered artificial neural networks Poirazi et al. (2003); Jadi et al. (2014); Ujfalussy et al. (2018); Tzilivaki et al. (2019); Beniaguev et al. (2021); Marino (2021); Jones and Kording (2021, 2022); Poirazi and Papoutsi (2020), and capable of discriminating between dozens to hundreds of input patterns Poirazi et al. (2003); Gütig and Sompolinsky (2006); Hawkins and Ahmad (2016); Moldwin and Segev (2020). Numerous biological mechanisms, such as complex ion channel dynamics (e.g. NMDA nonlinearity Major et al. (2013); Lafourcade et al. (2022); Tang et al. (2023)), plasticity on various and especially longer timescales (e.g. slow spike frequency adaptation Kobayashi et al. (2009); Bellec et al. (2020)), the intricate cell morphology (e.g. nonlinear integration by dendritic tree Stuart and Spruston (2015); Poirazi and Papoutsi (2020); Larkum (2022)), and their interactions, have been identified to contribute to their complexity.
Thus, one of the central questions in computational neuroscience is which learning and memory capabilities can be attributed to an individual cortical neuron and how mechanisms within a single neuron can work together to generate the observed computational complexity.
Detailed biophysical models of cortical neurons aim to capture this inherent complexity through high-fidelity simulations Hay et al. (2011); Herz et al. (2006); Almog and Korngreen (2016). However, they require a lot of computing resources to run and typically operate at a very fine level of granularity that does not facilitate the extraction of higher-level insights into the neuron’s computational principles. A promising approach to deriving such higher-level insights from simulations is through the training of surrogate phenomenological neuron models. These models are designed to replicate the output of high-fidelity biophysical simulations but use simplified interpretable components. This approach was employed, for example, to model computation in the dendritic tree via simple two-layer ANN Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Building on this line of research, a recent study by Beniaguev et al. (2021) developed a temporal convolutional network to capture the spike-level input/output (I/O) relationship with millisecond precision, accounting for the complexity of integrating diverse synaptic input across the entirety of the dendritic tree of a high-fidelity biophysical neuron model. It was found that a highly expressive temporal convolutional network with millions of parameters was essential to reproduce the aforementioned I/O relationship.
In this work, we propose that a model equipped with appropriate inductive biases that align with the operational principles of a cortical neuron should be capable of capturing the I/O relationship using a substantially smaller model size. To achieve this, a model would likely need to integrate multiple mechanisms of neural expressivity and judiciously allocate computational resources and parameters in close analogy to biological neurons. Should such a construction be possible, the requisite design choices may yield insights into higher-level principles of neural computation. We proceed to design the Expressive Leaky Memory (ELM) neuron model (see Figure 1), a biologically inspired, computationally expressive, yet parameter-efficient phenomenological model of a cortical neuron. The ELM neuron functions as a recurrent cell and can be conveniently used as a drop-in replacement for LSTMs Hochreiter and Schmidhuber (1997).
Our experiments show that a variant of the ELM neuron is expressive enough to accurately match the spike level I/O of a detailed biophysical model of a layer 5 pyramidal neuron at a millisecond temporal resolution with merely 8K parameters, in stark contrast to the millions of parameters required by temporal convolutional networks. To understand the implications of this design, we first probe the ELM neuron for its temporal information integration capabilities on a challenging biologically-inspired dataset requiring the addition of spike-encoded spoken digits. We find that the ELM neuron can outperform classic LSTMs, highlighting ELM neuron’s ability to leverage long-term memory. We subsequently evaluate the ELM neuron on three well-established long sequence modeling tasks prevalent in the machine learning literature, including the notoriously challenging Pathfinder-X task where it achieves $77.3\%$ accuracy even though many transformer-based models struggle to learn at all.
Our contributions are the following.
  1. 1.
    We propose the Expressive Leaky Memory (ELM) neuron model, a recurrent architecture that combines expressiveness with computational and parameter efficiency. The design of the ELM neuron is inspired by the structure and functionality of a biological cortical neuron.
  2. 2.
    Our experiments reveal that the ELM neuron effectively expresses the input/output relationship of a sophisticated biophysical model of a cortical neuron, suggesting that the inductive biases we introduced into the ELM neuron align with the operating principles of biological neurons.
  3. 3.
    We further demonstrate the ELM neuron’s ability to integrate temporal information over extended time horizons. Additionally, we show its capacity to handle complex long-range sequence modeling tasks that are common in machine learning, affirming its potential applicability in broader machine learning contexts.
2 The Expressive Leaky Memory Neuron
(a) A Cortical Neuron
Refer to caption (b) The ELM Neuron Architecture $\begin{split}{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\kappa_{s}}}&=\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\tau_{s}}}}})\\ {\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}&={\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\kappa_{s}}}\odot{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t-1}}}+{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}\odot\boldsymbol{x_{t}}\\ {\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}&=\tanh(\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}}([{\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}},{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}]))\\ {\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}&={\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}\odot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}+{\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}(1-{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}})\odot{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}\\ {\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{y_{t}}}&={\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}\cdot{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}\end{split}$ (1) (c) The ELM Neuron Equations
Figure 1: The biologically motivated Expressive Leaky Memory (ELM) neuron model. The architecture can be divided into the following components; the input current synapse dynamics, the integration mechanism dynamics, the leaky memory dynamics, and the output dynamics. a) Sketch of a biological cortical pyramidal neuron segmented into the analogous architectural components using the corresponding colors. b) Schematics of the ELM neuron architecture, component-wise colored accordingly. c) The ELM neuron equations, where $\boldsymbol{x_{t}}\in\mathbb{R}^{d_{s}}$ is the input at time $t$, $\Delta t\in\mathbb{R}^{+}$ the fictitious elapsed time in milliseconds between two consecutive inputs $\boldsymbol{x_{t-1}}$ and $\boldsymbol{x_{t}}$ , ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}\in\mathbb{R}^{d_{m}}$ are memory units, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s}}\in\mathbb{R}^{d_{s}}$ the synapse currents (traces), ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}\in\mathbb{R^{+}}^{d_{m}}$ and ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{\tau_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ their respective timescales in milliseconds, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}\in\mathbb{R^{+}}^{d_{s}}$ are synapse weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}$ the weights of a Multilayer Perceptron (MLP) with $l_{mlp}$ hidden layers of size $d_{\mathrm{mlp}}$, ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}\in\mathbb{R}^{d_{o}\times d_{m}}$ the output weights, ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}\in\mathbb{R^{+}}$ a scaling factor for the delta memory ${\color[rgb]{0.87109375,0.5625,0.01953125}\Delta m_{t}}\in\mathbb{R}^{d_{m}}$, and ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{y}}\in\mathbb{R}^{d_{o}}$ the output.
In this section, we discuss the design of the Expressive Leaky Memory (ELM) neuron, biologically motivating the individual computational components and the overall recurrent cell architecture.
2.1 Architectural Components
The current synapse dynamics.
Neurons typically receive inputs at their synapses in the form of sparse binary events, known as spikes, which are generally converted into small positive or negative currents exhibiting characteristic temporal dynamics Kandel et al. (2000). A prevalent abstraction in computational neuroscience for this dynamic is the current-based synapse model Dayan and Abbott (2005); Burkitt (2006), which we also adopt in the Expressive Leaky Memory (ELM) neuron model (refer to the green components in Figure 1). In essence, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}$ represents a low-pass filtered version of potentially sparse binary input, rendering it beneficial for coincidence detection or synaptic information integration König et al. (1996). The synapse weights, ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{w_{s}}}$, function as a straightforward gating mechanism, serving to either amplify or attenuate the input.
The memory unit dynamics.
The state of a biological neuron can be characterized by diverse measurable quantities, such as the membrane voltage, various ion or protein concentrations (e.g. Ca+, synaptic vesicles, ATP, etc.), and their slow or fast decay (determined by their respective timescale constants). Each of these quantities carries information through time, acting as a sort of memory mechanism or hidden state Kandel et al. (2000); Dayan and Abbott (2005) (see blue components in figure 1). Most computational neuroscience models a priori fix which quantities to track (usually the membrane voltage or certain ion concentrations and their interactions) and the relationships between them (through differential equations). However, which of these biological quantities are computationally relevant, how they interact, how many integration zones a neuron has, and what the relevant timescales of these processes remain topics of active debate Aru et al. (2020); Herz et al. (2006); Almog and Korngreen (2016); Koch (1997); Chavlis and Poirazi (2021). Therefore, to match a biological neuron’s computations and expressivity accurately, the surrogate model architecture needs to accommodate a large range of possibilities. In the ELM neuron, we achieve this by making the number of memory units $d_{m}$ a hyperparameter, unlike most other computational neuroscience models. Additionally, motivated by the large diversity of timescales found within and across neurons Cavanagh et al. (2020), and their computational importance Gjorgjieva et al. (2016); Gao et al. (2020); Zeraati et al. (2023), the memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t}}}$ are equipped with individually learnable memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, defining their respective rate of decay (large timescales meaning slower decay), constituting a learnable yet constant forgetting mechanism.
The integration mechanism dynamics.
This dynamic refers to how dependent on the previous memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}$ the synaptic input ${\color[rgb]{0.0078125,0.62109375,0.44921875}\boldsymbol{s_{t}}}$ is integrated into the memory unit updates ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$ in analogy to the dendritic tree of cortical neuron (orange components in Figure 1). We choose to parameterize this transformation using a Multilayer Perceptron (MLP), which is motivated by the ongoing discussion regarding the complexity and functional form of synaptic information integration over the dendritic tree in cortical neurons. While earlier perspectives suggested an integration process akin to linear summation Jolivet et al. (2008), newer studies advocate for complex non-linear integration Almog and Korngreen (2016); Gidon et al. (2020); Larkum (2022), specifically proposing multi-layered ANNs as a more suitable model Poirazi et al. (2003); Jadi et al. (2014); Tzilivaki et al. (2019); Marino (2021); Jones and Kording (2021, 2022); Hodassman et al. (2022), making an MLP a natural choice. Furthermore, its ability to incorporate previous memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m_{t-1}}}$ into the process accommodates state-dependent synaptic integration and related computations Hodgkin and Huxley (1952); Gasparini and Magee (2006); Bicknell and Häusser (2021), and enables the relationships between memory units ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}$ to be fully learnable, irrespective of their quantity. These features provide our neuron model with enhanced flexibility over those using pre-determined differential equations.
The output dynamics.
Spiking neurons emit their output spike at the axon hillock when the membrane voltage crosses an (approximate and varying) threshold Kandel et al. (2000). For the ELM neuron, we chose a flexible linear readout layer ${\color[rgb]{0.80078125,0.46875,0.73828125}\boldsymbol{w_{y}}}$ on top of the memory units (none of which have to correspond to membrane voltage a priori). The output of the readout layer can be a real value between 0 and 1, representing an output spike probability (in neuroscience attributed as a momentary rate), which ensures higher compatibility with typical machine learning tasks. Therefore, similar to biophysical models (e.g. Hodgkin Huxley Hodgkin and Huxley (1952)), there is no explicit reset mechanism, refractoriness, post-spike dynamics. All of these are implicitly learned by the ELM neuron. To instantiate an ELM neuron as a spiking neural network, a simple output rectification could be applied to recover binary output Zenke and Vogels (2021).
Additional design choices.
We found that enabling greater changes ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$ by introducing a multiplicative factor ${\color[rgb]{0.87109375,0.5625,0.01953125}\lambda}$ (a hyperparameter) improved training performance. Furthermore, note that the factor ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ is parametrized with $\exp(\nicefrac{{-\Delta t}}{{{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}}})$ (derived from classic neuron models Dayan and Abbott (2005)) and we learn the timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$ directly (constrained using a $sigmoid$ to the the desired range). As opposed to learning ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ and initializing it with $1-\Delta t/{\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, we get valid decay constants even for ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}<\Delta t$, and observed improved training stability.
3 Related Work
Accurately replicating the full spike-level input/output (I/O) relationship of detailed biophysical neuron models at millisecond resolution in a computationally efficient manner presents a formidable challenge. However, addressing this task could potentially yield valuable insights into neural mechanisms of expressivity, learning, and memory capabilities.
The relative scarcity of prior work on this subject can be partially attributed to the computational complexity of cortical neurons only recently garnering increased attention Tzilivaki et al. (2019); Beniaguev et al. (2021); Larkum (2022); Poirazi and Papoutsi (2020). Additionally, traditional phenomenological neuron models have primarily aimed to replicate specific computational phenomena of neurons or networks Koch (1997); Izhikevich (2004); Herz et al. (2006), rather than the entire I/O relationship. Finally, the computational resources and tooling necessary for an extensive exploration of suitable designs have only recently become widely available.
Phenomenological modeling research on temporally or spatially less detailed I/O relationship of biophysical neurons has been primarily centered around the use of multi-layered ANN structures in analogy to the neurons dendritic tree Poirazi et al. (2003); Tzilivaki et al. (2019); Ujfalussy et al. (2018). Similarly, we parametrize the synaptic integration with an MLP, while crucially extending this modeling perspective in several ways. Drawing upon the principles of classical phenomenological modeling via differential equations Izhikevich (2004); Dayan and Abbott (2005), our approach embraces the recurrent nature inherent to neurons. We further consider the significance of hidden states ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{m}}$ beyond membrane voltage, as seen in prior works with predetermined variables Brette and Gerstner (2005); Gerstner et al. (2014). This addition enables us to flexibly investigate internal memory timescales ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$, hinted at in earlier modeling studies Gjorgjieva et al. (2016); Cavanagh et al. (2020).
The relationship between timescales and the computations in the brain is the topic of numerous recent investigations in neuroscience. It was shown that intrinsic timescales are diverse across the cortex Murray et al. (2014), which can lead to increased neural expressivity Gjorgjieva et al. (2016); Cavanagh et al. (2020). A common way they occur in computational neuroscience is through the variables capturing slowly building up or leaking out biophysical quantities Hodgkin and Huxley (1952); Mozer (1991); Dayan and Abbott (2005). While timescales can be shaped by task requirements Gao et al. (2020); Zeraati et al. (2023); Cramer et al. (2020a), longer neuronal timescales might be used for temporal credit assignment across larger time horizons Bellec et al. (2018).
Deep learning architectures explicitly incorporating timescales have recently become more prevalent, with several recent architectures based on RNN, transformers or state-space models successfully incorporating them for improved temporal information processing Gu et al. ; Mahto et al. (2021); Smith et al. (2023); Ma et al. (2023). Initially, they appeared in Leaky-RNN Mozer (1991); Jaeger (2002); Kusupati et al. (2018); Tallec and Ollivier (2018), which use a convex combination of old memory and update (as done in ELM using ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$). In contrast to a time-varying memory decay or complex multi-staged gating architectures, the ELM features a much simpler recurrent cell architecture only using fixed but trainable time constants ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\tau_{m}}}$ and a single MLP to parametrize its integration dynamics.
Comparison with common RNN structures like LSTM Hochreiter and Schmidhuber (1997) and GRU Cho et al. (2014) surface two major differences. First, we set ${\color[rgb]{0.00390625,0.44921875,0.69921875}\boldsymbol{\kappa_{m}}}$ as a trainable constant instead of using parametric input and forget gating mechanisms Tallec and Ollivier (2018). Second, we utilize an MLP instead of a simpler parametric affine transformation for calculating ${\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{\Delta m_{t}}}$. Therefore, in contrast with most previous research, our architecture puts the primary importance not on the various gates, also leaving the forget gate constant during inference (which is otherwise believed to be hugely important Van Der Westhuizen and Lasenby (2018)), but rather on the memory update vector calculation itself.
4 Experiments
In the experimental section of this work, we address three primary research questions. First, are the inductive biases imparted in the ELM neuron adequate to accurately fit a high-fidelity biophysical simulation with a small number of parameters? We detail this investigation in Section 4.1. Second, is the ELM neuron capable of effectively integrating temporal information? We explore this issue in Section 4.2. Third, we test whether the ELM neuron is expressive enough to solve tasks with complex and long-range dependencies, which is discussed in Section 4.3. Each of these questions contributes to our overall understanding of the ELM neuron’s capabilities and its potential applications in both biological and machine learning contexts.
4.1 Fitting a complex biophysical cortical neuron model’s input-output relationship
The NeuronIO dataset primarily consists of simulated input-output (I/O) data for a complex biophysical layer 5 cortical pyramidal neuron model Hay et al. (2011). Input data features biologically inspired spiking patterns (1278 pre-synaptic spike channels), while output data comprises the model’s somatic membrane voltage and output spikes (see Figure 2a and 2b). The dataset and related code is publicly available Beniaguev et al. (2021). The dataset was pre-processed in accordance with Beniaguev et al. (2021), and the models were trained using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting.
Figure 2: The ELM neuron is a computationally efficient model of cortical neuron. a) detailed biophysical model of a layer 5 cortical pyramidal cell was used to generate the NeuronIO dataset consisting of input spikes and output spikes and voltage. b) Example target and predictions by the first ELM neuron that crosses the performance threshold. c) and d) Voltage and spike prediction performance of the respective surrogate models. While previous works required around 10M parameters to make accurate spike predictions using a TCN Beniaguev et al. (2021), an LSTM baseline is able to do it with 266K, and our ELM neuron model requires merely 53K, simultaneously achieving much better voltage prediction performance than the TCN. Additional figures comparing FLOPS in appendix.
Figure 3: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an ELM neuron with AUC $\approx 0.992$, and a (branch) ELM with the same default hyperparameters. Number of removed divergent runs marked with $1^{*}$. a) We find between 10 and 20 memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich (2004); Dayan and Abbott (2005). b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart and Spruston (2015); Jones and Kording (2022); Larkum (2022). c) Allowing greater updates to the memory units is beneficial. d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with default range being 1ms-150ms. Timescales around 25 ms seem to be the most useful (matching the typical membrane timescale in the cortex Dayan and Abbott (2005)); however, a lack can be partially compensated by longer timescales. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$ of the (branch) ELM neuron.
Our ELM neuron achieves better prediction of voltage and spikes than previously used architectures for any given number of trainable parameters (and compute). In particular, it crosses the “sufficiently good” spike prediction performance threshold (0.991 AUC) as defined in (Beniaguev et al., 2021) by using merely 50K trainable parameters, which is around 200$\times$ improvement compared to the previous attempt (TCN) that required around 10M trainable parameters, and 6$\times$ improvement over a LSTM baseline which requires around 266K parameters (see Figure 2c-d). Overall, this result indicates that recurrent computation is indeed an appropriate inductive bias for modeling cortical neurons.
We use the fitted model to investigate how many memory units and which timescales are needed to match the neuron closely. We find that around 20 memory units are required (Figure 3a) with timescales that are allowed to reach at least 25 ms (Figure 3d). While a diversity of timescales including long ones seem to be favorable for accurate modeling (Figure 3d and 3f), using constant memory timescales around 25 ms performs surprisingly well (matching the typical membrane timescales used in computational modeling Dayan and Abbott (2005), Figure 3e). Removing the hidden layer or decreasing the integration mechanism complexity deteriorates the performance significantly (Figure 3b). Allowing for more rapid changes of the memory through larger $\lambda$ is necessary (Figure 3c), possibly to match the fast internal dynamics of neurons around spike times.
How much nonlinearity is in the dendritic tree?
Within the ELM architecture, we allow for nonlinear interaction between any two synaptic inputs via the MLP. This flexibility might be necessary in cases where little is a priori known about the structure of the input. However, for matching the I/O of biological neurons, knowledge of neuronal morphology and biophysical assumptions about linear-nonlinear computations in the dendritic tree might be exploited to reduce the dimensionality of the input to the MLP, which so far is particularly costly due to the large number of synaptic inputs it receives ($d_{s}=1278$). In particular, many studies suggest individual dendritic branches compute a rectified sum of their synaptic inputs Poirazi et al. (2003); Stuart and Spruston (2015); Hawkins and Ahmad (2016); Poirazi and Papoutsi (2020), with additional nonlinearities located downstream. Consequently, we modify the ELM neuron to include virtual branches along which the synaptic input is first reduced by a simple summation before further processing (see Figure 4. For NeuronIO specifically, we assign the synaptic inputs to the branches in a moving window fashion (exploiting that in the dataset, neighboring inputs were also typically neighboring synaptic contacts on the same dendritic branch of the biophysical model). The window size is controlled by the branch size, and number of branches implicitly defines the stride size to ensure equally spaced sampling across the $d_{s}=1278$ inputs.
Figure 4: Coarse-grained modeling of synaptic integration significantly improves model efficiency. a) The integration mechanism dynamics of the ELM now computes the activity of individual dendritic branches as a simple sum of their respective synaptic inputs first before passing them on to the $\text{MLP}_{{\color[rgb]{0.87109375,0.5625,0.01953125}\boldsymbol{w_{p}}}}$, where $d_{\mathrm{tree}}$ is the number of branches and $d_{\mathrm{brch}}$ the number of synapses per branch. b-c) Produced using joint ablation of $d_{m}$ with $d_{\mathrm{mlp}}=2d_{m}$. The new (branch) ELM neuron further improves on the vanilla ELM neuron by about 7$\times$ in terms of parameter efficacy. A (branch) ELM model with a humble 8104 trainable parameters now reaches 0.9924 AUC (3rd from left), using otherwise same hyper-parameters as the ELM (6th from right).
Surprisingly, even with this strong simplification, the (branch) ELM neuron model is able to retain its predictive performance while requiring a modest 8K trainable parameters to substantially cross the performance threshold. This corresponds to roughly 7$\times$ reduction over the vanilla ELM neuron. We also find that a combination of $d_{\mathrm{tree}}=35$ and $d_{\mathrm{brch}}=65$ still achieved over $0.991$ AUC with only 5479 trainable parameters, corroborating the assumption of the near-linear computation within dendritic branches and inviting future investigation of minimal required synaptic nonlinearity. However, this simplification ideally utilizes the knowledge of the exact morphology for modeling the neuron, or the relationship structure in the input data in case of other tasks; thus most consecutive experimental reports are for the vanilla ELM neuron.
4.2 Evaluating temporal information integration capabilities on a bio-inspired task
The Spiking Heidelberg Digits (SHD) dataset comprises spike-encoded spoken digits (0-9) in German and English Cramer et al. (2020b). The digits were encoded using 700 input channels in a biologically-inspired artificial cochlea. Each channel represents a narrow frequency band with the firing rate coding for the signal power in this band, resulting in an encoding that resembles the spectrogram of the spoken digit (see Figure 5a).
Motivated by recent findings that most neuromorphic benchmark datasets only require very limited temporal processing abilities Yang et al. (2021), we introduce the SHD-Adding dataset, created by concatenating two uniformly and independently sampled SHD digits and setting the target to their sum (regardless of language) (see Figure 5a). Solving this dataset necessitates identifying each digit on a shorter timescale and computing their sum by integrating this information over a longer timescale, which in turn requires retaining the first digit in memory. Whether single cortical neurons can solve this exact task is unclear; however, it has been shown that even single neurons possibly encode and perform basic arithmetics in the medial temporal lobe Cantlon and Brannon (2007); Kutter et al. (2018, 2022).
Figure 5: The ELM neuron performs well on long and sparse data using longer timescales. a) Sample from the biologically motivated SHD-Adding dataset (based on Cramer et al. (2020b)), each dot is an input spike, and a vertical dashed line is a guide for the eye indicating the separation of the two digits (not communicated to the network). b-d) The ELM neuron consistently outperforms a classic LSTM, especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. Our Leaky Integrate and Fire (LIF) Spiking Neural Network (SNN) does not manage to achieve good performance for any bin size, and training becomes unstable for long sequences. e) and f) Ablations using bin size of 2ms with test set performance reported. e) Solving SHD-Adding requires ELM neuron to have a higher complexity than required for NeuronIO. f) Longer memory timescales are crucial for extracting long-range dependencies in the data.
The ELM neuron demonstrates its ability to solve the summing task across various temporal resolutions, as determined by the bin size. As we vary the bin size from 1ms (comprising 2000 bins in total, corresponding to the maximum temporal detail and longest required memory retention) to 100 (with 20 bins in total, corresponding to the minimum temporal detail and shortest memory retention), the ELM neuron’s performance remains robust, degrading only gracefully for greater bin sizes (see Figure 5b-d). Further, the performance is also maintained when testing on two held-out speakers, showing that the ELM neuron remains comparatively robust out-of-distribution. In contrast, the LSTM struggles with this task, especially when the bin size is below 50, due to memory issues. As the bin size increases, the LSTM’s performance does not improve because larger bin sizes leads to the loss of crucial temporal details. This outcome underlines the importance of a model’s ability to integrate complex synaptic information effectively (see Figure 5e). It also highlights the need for models to accommodate longer memory timescales, in light of the observation that limiting our model to memory timescales below 150 has a severe impact on performance (see Figure 5f).
4.3 Evaluating on complex and very long temporal dependency tasks
To test the extent of the ELM neuron’s ability to extract complex long-range dependencies, we use three modern machine learning tasks. The Sequential CIFAR-10 task involves solving the 10-way CIFAR-10 Krizhevsky et al. (2009) classification problem given a flattened sequence of $1024$ pixel values, corresponding to $32\times 32$ sized images Tay et al. (2021). The Pathfinder and Pathfinder-X tasks Linsley et al. (2018); Tay et al. (2021) require determining whether two dots are connected by a path in a 2D image (see Figure 6, but the images are flattened out to be a sequence of pixels before being fed into the model. For the Pathfinder-X task, the length of this sequence is $16384$, corresponding to a $128\times 128$ image (see Figure 6), making it a notoriously difficult long-range modeling task.
Our results are summarized in Table 1, where we compare the ELM neuron against several strong baselines. The model most comparable to ours is an LSTM with explicit long timescale-derived gating bias initialization Tallec and Ollivier (2018) (Chrono-LSTM). We also incorporate two self-attention-based models, including the Transformer Vaswani et al. (2017) and the Longformer Bellec et al. (2018). Finally, we include Structured State Space Sequence model Gu et al. (S4) and a Transformer leveraging exponential moving averages Ma et al. (2023) (Mega), the latter being the current state-of-the-art.
When comparing between recurrent cells (ELM Neuron vs. Chrono-LSTM), we find that both models solve the tasks, albeit the ELM Neuron substantially outperforms Chrono-LSTM on the Pathfinder-X task. The self-attention-based models trail further behind, with both Transformer and Longformer failing to solve the Pathfinder-X task Tay et al. (2021). S4 and Mega perform better, but they are both deep architectures; Mega features 6 layers of 128 and 256 memory cells and self-attention, compared to the ELM neuron with 150 memory units and less than $100$K parameters.
Table 1: Long Sequence Modeling Datasets
Model Accuracy
Sequential
CIFAR-10
Pathfinder Pathfinder-X
ELM Neuron (ours) 54.0 74.72 77.3
Chrono-LSTM Tallec and Ollivier (2018) 46.61 70.79 51.85
Transformer Vaswani et al. (2017) 42.44 71.4 FAIL
Longformer Bellec et al. (2018) 42.22 69.71 FAIL
S4 Gu et al. 87.26 86.05 88.1
Mega Ma et al. (2023) 90.44 96.01 97.98
The ELM neuron can solve long-range sequence modeling tasks. Results on three complex temporal dependency tasks. Surprisingly, the simple ELM architecture is able to solve the notoriously challenging Pathfinder-X task with sample length $>16$K. It scores higher than the transformer baselines and is only outperformed by big 6-staged architectures tuned specifically for these tasks (S4, Mega).
[Uncaptioned image] Figure 6: Pathfinder-X Samples. The task is to classify whether two dots are connected by a dashed line. Each $128\times 128$ image is flattened to a sequence in raster-scan order, yielding sequences of total length $16384$.
Overall, the results suggest that the ELM neuron is capable of solving tasks with very long temporal dependencies and doing so without employing multistage structures. We hypothesize that assembling the ELM neurons into layered network structures might give it enough processing capabilities to catch up with the deep models but leave this investigation to future work.
5 Conclusion
In this study, we introduced a biologically inspired recurrent cell, the Expressive Leaky Memory (ELM) neuron, and demonstrated its capability to fit the full spike-level input/output mapping of a high-fidelity biophysical neuron model. Unlike previous works that achieved this fit with millions of parameters, our a variant of our model only requires eight thousand, thanks to the careful design of the architecture leveraging appropriate implicit biases. Furthermore, the ELM neuron can effectively model a cortical neuron without making rigid assumptions about the number of memory states, unlike existing phenomenological models. It also imposes fewer presumptions about the complexity of the integration mechanism and the exact values of the timescales associated with each memory state.
We further scrutinized the implications of this design on various tasks, including some notoriously challenging ones in the machine learning literature. The ELM neuron was found to be competitive, a notable feat considering its biological motivation, a characteristic often lacking in similar successful models. However, we acknowledge the ongoing debate regarding the degree of "biological plausibility" required in intelligent systems. Biological systems, including neurons, are subject to a variety of ’wetware’ constraints that may not need to be reproduced.
Despite this caveat, the promising performance of the ELM neuron on machine learning tasks suggests that we are beginning to incorporate the inductive biases that drive the development of more intelligent systems. Future research focused on connecting ELM neurons into a network could provide even more insights into the necessary and dispensable elements for building smarter machines.
Acknowledgments and Disclosure of Funding
This work was supported by a Sofja Kovalevskaja Award from the Alexander von Humboldt Foundation. We acknowledge the support from the BMBF through the Tübingen AI Center (FKZ: 01IS18039A and 01IS18039B). AL, GM, and BS are members of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 39072764.
Appendix
Appendix A Implementation Details
All computations were performed using Python 3.9, and the following libraries were instrumental in our implementation: jax 0.3.14 (coupled with jaxlib 0.3.10 as a GPU back-end) for auto-grad and auto-vectorization; equinox 0.8.0, a jax-based neural network library; optax 0.1.3, a jax-based optimizer library; and pytorch 1.12.1 for data-loading. The code used for our experiments will be accessible at the public repository upon publication. To maintain anonymity, the essential excerpts of the code are presented in Section E.
Table S1: The default ELM neuron parameters and recommendations.
Hyper-Parameter NeuronIO Tuning Recommendation
$\boldsymbol{s_{0}}$ 0 /
$\boldsymbol{w_{s}}$ 0.5 /
$\boldsymbol{w_{s}}$ (bounds) > 0 >= 0
$\boldsymbol{\tau_{s}}$ 5ms greater for long and sparse data
$\boldsymbol{m_{0}}$ 0 /
$d_{m}$ $20$ between $10-250$
$\boldsymbol{\tau_{m}}$ (init technique) equally spaced evenly spaced on logarithmic scale
$\boldsymbol{\tau_{m}}$ (init range) 1ms,100ms 1ms to sample length in ms
$\boldsymbol{\tau_{m}}$ (bounds) 0ms,500ms same as init range or larger
$\text{MLP}_{w_{p}}$ (nonlineary) ReLU /
$\text{MLP}_{w_{p}}$ (bias) True /
$\text{MLP}_{w_{p}}$ (init technique) Kaiming Uniform /
$l_{\mathrm{mlp}}$ $1$ /
$d_{\mathrm{mlp}}$ $2*{\color[rgb]{0.00390625,0.44921875,0.69921875}d_{m}}$ $2*{\color[rgb]{0.00390625,0.44921875,0.69921875}d_{m}}$
$\lambda$ 10 between $1-10$
$\boldsymbol{w_{y}}$ (bias) True /
$\boldsymbol{w_{y}}$ (init technique) Kaiming Uniform /
$d_{\mathrm{tree}}$ 45 receive all input 2-5 times over
$d_{\mathrm{brch}}$ 100 $2*{\color[rgb]{0.3359375,0.70703125,0.9140625}d_{\mathrm{tree}}}$
Recommended default and tuning parameters:
On a novel dataset, we recommend starting with a smaller $d_{m}$ around $100$, and small $\lambda$ around $3$, as increasing both simultaneously may lead to increased training instability. Potential future work could explore setting $\lambda$ automatically based on $d_{m}$, or learning it directly. The timescales $\boldsymbol{\tau_{m}}$ should generally be derived from the dataset length and the suspected timescales of the temporal dependencies within the data. Increasing $\boldsymbol{\tau_{s}}$ may help to enhance learning speed in some cases. When using the (branch) ELM it is important to sufficiently over-sample the input (more synapses than input) as significant expressivity stems from $\boldsymbol{w_{s}}$ doing the selection. Additional recommendations are summarized in Table S1.
Appendix B Datasets and Training Details
General training setup:
For each task and dataset, the training dataset was deterministically split to create a consistent validation dataset, which was used for model selection during training and hyperparameter tuning. By default all models were trained using Backpropagation Through Time (BPTT) with a batch size of 8 and the Adam optimizer, employing a learning rate of $5e^{-4}$ and a cosine-decay schedule across the entire training duration. All experiments were run on a single A100-40GB or A100-80GB and ran less than 24h, Pathfinder-X being the notable exception.
NeuronIO Dataset:
For training and evaluation the dataset was pre-processed in accordance with Beniaguev et al. [2021], by capping somatic membrane voltage at -55mV and subtracting a bias of -67.7mV. Additionally, the somatic membrane voltage was scaled by 1/10 for training. Training samples were 500ms long with a 1ms bin size and $\Delta t$. The ELM neuron used the default parameters from Table S1. Models were trained for 30 epochs of 11,400 batches per epoch using Binary Cross Entropy (BCE) for spike prediction and Mean Squared Error (MSE) for somatic voltage prediction, with equal weighting. Loss was calculated after a 150ms burn-in period. The mean and standard deviation over three runs is reported, with Root Means Squared Error (RMSE) and Area Under the Receiver Operator Curve (AUC) for voltage and spike prediction, respectively.
Spiking Heidelberg Digits (Adding) Datasets:
The digits were preprocessed by cutting them to a uniform length of one second and binning the spikes using various bin sizes, the default being 10ms. The models were trained using the Adamax optimizer with a learning rate of $5e^{-3}$ for 70 epochs, for 814 and 2000 batches per epoch for SHD and SHD-Adding respectively, with $\Delta t$ set to the bin size, and dropout probability set to 0.5. The ELM used $\lambda=5$, $d_{m}=100$ and $\tau_{m}$ initialized evenly spaced between 1ms and 150ms with bounds of 0ms to 1000ms, and the LSTM used a hidden size of 250 and additional recurrent dropout of 0.3, while the SNN used a learning rate of $2e^{-3}$ no dropout but a $l1$ regularization on the spikes of $0.01$. Models were trained using the Cross-Entropy (CE) loss on the last float output of the respective model, and the performance was reported as prediction Accuracy, with mean and standard deviation calculated over five runs (chance performance being 1/19).
Long Sequence Modeling Datasets: The images were preprocessed by binning the 256 greyscale values into 16 values, and 32 different values for Pathfinder-X. The models were trained for 300 epochs with a batch size of 384, and batch size of 768 for Pathfinder-X. All trained using Cross-Entropy (CE) loss on the last output of the model, and the performance is reported as prediction Accuracy. The mean over three runs is reported for all experiments.
Table S2: The ELM neuron configuration
Dataset Data Bins Batch Size LR $d_{m}$ $\lambda$ Timescales
CIFAR-10 16 384 $1e^{-3}$ 150 3 logspace: $1-10^{3}$
Pathfinder 16 384 $3e^{-4}$ 150 3 logspace: $1-10^{3}$
Pathfinder-X 32 768 $2e^{-4}$ 150 5 logspace: $1-10^{4}$
The ELM memory timescale bounds were matched to the initialization range. On the Pathfinder dataset an additional input dropout of $0.2$ was applied, however using $lr=2e^{-4}$ and no dropout still achieves $71.61\%$ accuracy. Furthermore, the ELM used a synapse tau of $150ms$ on the Pathfinder-X dataset, which we observed to increase the learning speed.
Table S3: The Chrono-LSTM configuration
Dataset Data Bins Batch Size LR Hidden Size Timescales
CIFAR-10 $16$ $384$ $1e^{-3}$ $150$ uniform: $1-2*10^{3}$
Pathfinder $16$ $384$ $1e^{-3}$ $150$ uniform: $1-2*10^{3}$
Pathfinder-X $32$ $768$ $1e^{-3}$ $70$ uniform: $1-2*10^{4}$
Alternative configurations for the Chrono-LSTM model, including learning rates of $2e^{-4}$, $5e^{-4}$, and $7e^{-4}$, as well as a hidden size of $150$ with half the batch size (reaching the 80GB GPU memory limit), were evaluated. For Pathfinder-X, only the reported configuration demonstrated performance exceeding chance levels, and only for one of the three trials.
Appendix C Additional Results
Figure S1: The ELM neuron is a computationally efficient model of cortical neuron. Similar figure to 2c and 2d, however, displaying FLOPs required to do inference on a single sample. a) and b) Voltage and spike prediction performance of the respective surrogate models. While previous works required around 10M parameters to make accurate spike predictions using a TCN Beniaguev et al. [2021], an LSTM baseline is able to do it with 266K, our ELM neuron model requires merely 53K, and our (branch) ELM neuron only a humble 8K, simultaneously achieving much better voltage prediction performance than the TCN. A throughput optimized ELM neuron implementation can potentially reduce the required FLOPs even further.
Figure S2: The ELM neuron gives relevant neuroscientific insights. Ablations on NeuronIO of different hyperparameters of an (branch) ELM neuron with AUC $\approx 0.992$, with default hyperparameters. The number of removed divergent runs marked with $1^{*}$. a) We find above memory-like hidden states to be required for accurate predictions, much more than typical phenomenological models use Izhikevich [2004], Dayan and Abbott [2005]. b) Highly nonlinear integration of synaptic input is required, in line with recent neuroscientific findings Stuart and Spruston [2015], Jones and Kording [2022], Larkum [2022]. c) Allowing greater updates to the memory units is beneficial, however, too large ones increase training instability. d-f) Ablations of memory timescale (initialization and bounding) range or (constant) value, with the default range being 1ms-150ms. Timescales around 25ms-50ms seem to be the most useful (matching the typical membrane timescale in the cortex Dayan and Abbott [2005]); however, a lack can be partially compensated by longer timescales, even better than by the vanilla ELM. g) and h) Ablating the number of branches $d_{\mathrm{tree}}$ and number of synapses per branch $d_{\mathrm{brch}}$.
Figure S3: The ELM neuron performs well on typical spiking datasets. The following results are on the original Spiking Heidelberg Digits dataset Cramer et al. [2020b]. a-c) The ELM neuron consistently outperforms a classic LSTM, especially for smaller bin sizes (meaning longer training samples), and LSTM-performance cannot be fully recovered even for larger bin sizes. Our Leaky Integrate and Fire (LIF) Spiking Neural Network (SNN), however, does manage to achieve decent performance for intermediate bin sizes around 20, in contrast to the SHD-Adding dataset (see Figure 5).
Appendix D Additional Visualizations
Figure S4: Visualization of ELM neuron dynamics. Extended visualization of Figure 2b for an ELM neuron achieving around 0.992 AUC. a) The synaptic input as the neuron model receives it, with excitatory input ($+1$) marked in red, and inhibitory input ($-1$) marked in blue. b) The ELM neurons predictions and the ground-truth targets, for a particularly challenging example. c) A random subset of 20 synapses are visualized. Synapses receiving negative input will be deflected downwards. d) All 20 memory values are visualized. Some fluctuate more rapidly than others, typically proportional to their memory timescales.
Appendix E Reduced Implementation In Pytorch
import math
import torch
import torch.nn as nn
# NOTE: the 1e-6 used throughout are to avoid numerical instabilities
# NOTE: the tau_m bounds are enforced using a scaled sigmoid
# NOTE: the w_s bounds are enforced using relu (originally exp)
def scaled_sigmoid(x, lower_bound, upper_bound):
    return (upper_bound - lower_bound) * torch.sigmoid(x) + lower_bound
def inverse_scaled_sigmoid(x, lower_bound, upper_bound):
    x = torch.clamp(x, lower_bound + 1e-6, upper_bound - 1e-6)
    return torch.log((x - lower_bound) / (upper_bound - x))
def custom_tanh(x):
    return torch.tanh(x * 2 / 3) * 1.7159
class MLP(nn.Module):
    """Multi-layerperceptron(MLP)model."""
    def __init__(self, input_size, hidden_size, num_output, num_layers):
        super(MLP, self).__init__()
        next_input_size = input_size
        layers = []
        for i in range(num_layers):
            layers.append(nn.Linear(next_input_size, hidden_size))
            layers.append(nn.ReLU())
            next_input_size = hidden_size
        layers.append(nn.Linear(next_input_size, num_output))
        self.network = nn.Sequential(*layers)
    def forward(self, x):
        return self.network(x)
class ELM(nn.Module):
    """ExpressiveLeakyMemory(ELM)neuronmodel."""
    def __init__(self, num_synapse, num_output, num_memory=100,
                 mlp_num_layers=1, memory_tau_min=10, memory_tau_max=1000,
                 lambda_value=3.0, tau_s_value=5.0, w_s_value=0.5, delta_t=1.0,
                 learn_memory_tau=False,mlp_hidden_size=None,use_num_branch=None):
        super(ELM, self).__init__()
        self.delta_t = delta_t
        self.lambda_value = lambda_value
        self.learn_memory_tau = learn_memory_tau
        self.memory_tau_min = memory_tau_min
        self.memory_tau_max = memory_tau_max
        # configuration in case of using branches
        self.use_num_branch = use_num_branch
        if use_num_branch is not None:
            assert num_synapse % use_num_branch == 0, "mustbeexactlydivisible"
            num_mlp_input = use_num_branch + num_memory
        else:
            num_mlp_input = num_synapse + num_memory
        # initialization of model parameters
        mlp_hidden_size = mlp_hidden_size if mlp_hidden_size else 2 * num_memory
        self.mlp = MLP(num_mlp_input, mlp_hidden_size, num_memory, mlp_num_layers)
        self.w_s = nn.parameter.Parameter(torch.full((num_synapse,), w_s_value))
        self.w_y = nn.Linear(num_memory, num_output)
        # initialization of time constants and decay factors
        self.tau_s = torch.full((num_synapse,), tau_s_value)
        self.kappa_s = torch.exp(-delta_t / torch.clamp(self.tau_s, min=1e-6))
        self._proto_tau_m = torch.logspace(
            math.log10(memory_tau_min + 1e-6),
            math.log10(memory_tau_max - 1e-6),
            num_memory,
        self._proto_kappa_m = torch.exp(
            -delta_t / torch.clamp(self._proto_tau_m, min=1e-6)
        # configuration in case memory tau is learnable
        if self.learn_memory_tau:
            self._proto_tau_m = inverse_scaled_sigmoid(
                self._proto_tau_m, memory_tau_min, memory_tau_max
            self._proto_tau_m = nn.parameter.Parameter(self._proto_tau_m)
    @property
    def tau_m(self):
        # ensure that tau_m is always up to date
        if self.learn_memory_tau:
            return scaled_sigmoid(
                self._proto_tau_m, self.memory_tau_min, self.memory_tau_max
        else:
            return self._proto_tau_m
    @property
    def kappa_m(self):
        # ensure that kappa_m is always up to date
        if self.learn_memory_tau:
            return torch.exp(-self.delta_t / torch.clamp(self.tau_m, min=1e-6))
        else:
            return self._proto_kappa_m
    def dynamics(self, x, s_prev, m_prev, kappa_m):
        # compute the dynamics for a single timestep
        batch_size = x.shape[0]
        s_t = self.kappa_s * s_prev + torch.relu(self.w_s) * x
        if self.use_num_branch is not None:
            syn_input = s_t.view(batch_size, self.use_num_branch, -1).sum(dim=-1)
        else:
            syn_input = s_t
        delta_m_t = custom_tanh(
            self.mlp(torch.cat([syn_input, kappa_m * m_prev], dim=-1))
        m_t = kappa_m * m_prev + self.lambda_value * (1 - kappa_m) * delta_m_t
        y_t = self.w_y(m_t)
        return y_t, s_t, m_t
    def forward(self, X):
        # compute the the recurrent dynamics for a sample
        batch_size, T, num_synapse = X.shape
        kappa_m = self.kappa_m
        s_prev = torch.zeros(batch_size, len(self.tau_s), device=X.device)
        m_prev = torch.zeros(batch_size, len(self.tau_m), device=X.device)
        outputs = []
        for t in range(T):
            y_t, s_prev, m_prev = self.dynamics(X[:, t], s_prev, m_prev, kappa_m)
            outputs.append(y_t)
        return torch.stack(outputs, dim=-2)
Appendix F Results in Table Format
Table S4: The (branch) ELM on NeuronIO
Trainable
Parameters
FLOPs
Voltage Prediction
RMSE
Spike Prediction
AUC
Memory Units
10 5.9K 16.04M 0.695 ± 0.004 0.989 ± 0.0001
15 6.9K 17.06M 0.652 ± 0.0 0.9913 ± 0.0007
20 8.1K 18.27M 0.638 ± 0.003 0.9924 ± 0.0002
25 9.5K 19.69M 0.639 ± 0.013 0.9922 ± 0.0006
30 11.1K 21.3M 0.628 ± 0.002 0.9927 ± 0.0001
40 14.9K 25.13M 0.633 ± 0.014 0.9925 ± 0.0005
50 19.5K 29.76M 0.629 ± 0.012 0.9925 ± 0.0009
Table S5: The ELM on NeuronIO
Trainable
Parameters
FLOPs
Voltage Prediction
RMSE
Spike Prediction
AUC
Memory Units
10 26.06K 28.49M 0.709 ± 0.001 0.9878 ± 0.0004
15 39.39K 41.84M 0.662 ± 0.003 0.99 ± 0.0013
20 52.92K 55.38M 0.643 ± 0.004 0.9918 ± 0.0002
25 66.65K 69.13M 0.648 ± 0.008 0.9918 ± 0.0002
35 94.71K 97.22M 0.631 ± 0.001 0.9922 ± 0.0001
50 138.3K 140.85M 0.629 ± 0.002 0.9925 ± 0.0002
75 214.95K 217.58M 0.622 ± 0.004 0.9927 ± 0.0001
100 296.6K 299.3M 0.623 ± 0.006 0.9927 ± 0.0001
Table S6: The LSTM on NeuronIO
Trainable
Parameters
FLOPs
Voltage Prediction
RMSE
Spike Prediction
AUC
Hidden Size
15 77.67K 77.77M 0.717 ± 0.021 0.9864 ± 0.0008
25 130.45K 130.61M 0.673 ± 0.003 0.9898 ± 0.0002
50 265.9K 266.23M 0.641 ± 0.004 0.9913 ± 0.0002
100 551.8K 552.45M 0.624 ± 0.002 0.9923 ± 0.0002
150 857.7K 858.68M 0.623 ± 0.005 0.9925 ± 0.0001
250 1529.5K 1531.13M 0.626 ± 0.005 0.9925 ± 0.0004
Table S7: SHD Results
Train Accuracy Valid Accuracy Test Accuracy
Model Bin Size
ELM 1 1.0 ± 0.0 0.98 ± 0.0 0.9 ± 0.0
2 1.0 ± 0.0 0.98 ± 0.0 0.91 ± 0.01
4 1.0 ± 0.0 0.98 ± 0.0 0.9 ± 0.01
6 1.0 ± 0.0 0.98 ± 0.0 0.9 ± 0.02
8 1.0 ± 0.0 0.98 ± 0.0 0.89 ± 0.01
10 1.0 ± 0.0 0.98 ± 0.0 0.89 ± 0.0
20 1.0 ± 0.0 0.97 ± 0.0 0.88 ± 0.01
50 0.99 ± 0.01 0.95 ± 0.01 0.84 ± 0.01
100 1.0 ± 0.0 0.93 ± 0.01 0.77 ± 0.01
LSTM 1 0.26 ± 0.41 0.24 ± 0.4 0.21 ± 0.35
2 0.29 ± 0.34 0.27 ± 0.33 0.21 ± 0.3
4 0.94 ± 0.07 0.91 ± 0.07 0.78 ± 0.09
6 0.99 ± 0.0 0.96 ± 0.01 0.83 ± 0.02
8 1.0 ± 0.0 0.96 ± 0.0 0.86 ± 0.03
10 1.0 ± 0.0 0.97 ± 0.0 0.87 ± 0.01
20 1.0 ± 0.0 0.96 ± 0.0 0.82 ± 0.01
50 1.0 ± 0.0 0.94 ± 0.01 0.81 ± 0.0
100 0.99 ± 0.0 0.91 ± 0.01 0.74 ± 0.0
SNN 1 0.05 ± 0.0 0.05 ± 0.0 0.05 ± 0.0
2 0.07 ± 0.02 0.07 ± 0.02 0.07 ± 0.02
4 0.16 ± 0.04 0.16 ± 0.05 0.16 ± 0.05
6 0.36 ± 0.04 0.35 ± 0.05 0.35 ± 0.05
8 0.46 ± 0.04 0.45 ± 0.04 0.45 ± 0.04
10 0.51 ± 0.07 0.49 ± 0.07 0.5 ± 0.06
20 0.61 ± 0.04 0.59 ± 0.05 0.58 ± 0.04
50 0.49 ± 0.02 0.47 ± 0.02 0.47 ± 0.02
100 0.41 ± 0.03 0.39 ± 0.02 0.39 ± 0.03
Table S8: SHD-Adding Results
Train Accuracy Valid Accuracy Test Accuracy
Model Bin Size
ELM 1 1.0 ± 0.0 0.96 ± 0.01 0.82 ± 0.01
2 1.0 ± 0.0 0.96 ± 0.0 0.82 ± 0.01
4 0.98 ± 0.03 0.93 ± 0.03 0.81 ± 0.03
6 0.95 ± 0.05 0.89 ± 0.05 0.76 ± 0.07
8 0.94 ± 0.09 0.88 ± 0.09 0.76 ± 0.08
10 0.98 ± 0.04 0.93 ± 0.04 0.8 ± 0.03
20 0.99 ± 0.0 0.94 ± 0.01 0.79 ± 0.02
50 0.99 ± 0.0 0.92 ± 0.01 0.75 ± 0.01
100 0.98 ± 0.0 0.86 ± 0.01 0.66 ± 0.02
LSTM 1 0.1 ± 0.0 0.1 ± 0.0 0.1 ± 0.01
2 0.1 ± 0.0 0.1 ± 0.0 0.1 ± 0.0
4 0.1 ± 0.0 0.1 ± 0.0 0.1 ± 0.0
6 0.1 ± 0.0 0.1 ± 0.0 0.1 ± 0.0
8 0.11 ± 0.02 0.11 ± 0.02 0.1 ± 0.01
10 0.21 ± 0.08 0.19 ± 0.07 0.15 ± 0.05
20 0.83 ± 0.37 0.75 ± 0.33 0.6 ± 0.26
50 0.99 ± 0.0 0.87 ± 0.01 0.67 ± 0.02
100 0.98 ± 0.0 0.75 ± 0.01 0.55 ± 0.01
SNN 10 0.09 ± 0.01 0.09 ± 0.01 0.08 ± 0.01
20 0.08 ± 0.01 0.08 ± 0.01 0.08 ± 0.01
50 0.1 ± 0.01 0.1 ± 0.01 0.1 ± 0.0
100 0.09 ± 0.0 0.09 ± 0.0 0.09 ± 0.0